show you work pls
4. Complete the Dupont Chart. 2017 2018 2019 ROE 18.20% 19.25% EM 1.25 1.33 ROA 13.80% 13.23% PM 10.00% 11.00% TAT 1.5
Answer #1

First let's understand Dupont Analysis.

Return on Equity is the return generated as a percentage of total equity, so ROE = Net Income / Total Equity

To calculate ROE using other ratios, the formula is expanded by multiplying and dividing the ratio with sales and total assets, So the equation becomes ROE = (Net Income/Sales) * (Sales/Total Assets) * (Total Assets/Total Equity)

ROE = Profit margin (PM) * Total Asset Multiplier (TAT) * Equity multiplier (EM) = PM*TAT*EM

Return on Assets (ROA) = Net Income / Total Assets, Expanding this in similar fashion gives

ROA = (Net Income/Sales) * (Sales/Total Assets) = PM * TAT

In the above table, we can fill in the blanks using these 2 formulae

Given Table

2017 2018 2019
ROE 18.20% 19.25% E
EM 1.25 C 1.33
ROA A 13.80% 13.23%
PM 10.00% D 11.00%
TAT B 1.50 F

ROE = PM * TAT * EM ; 18.20% = 10%*B*1.25 ; B = 1820 / 125 = 1.456 ~ 1.46

ROA = PM * TAT ; A = 10% * 1.46 = 14.60%

ROA = PM*TAT ; 13.80% = D * 1.5 ; D = 9.20%

ROE = PM*TAT*EM ; 19.25% = 9.20% * 1.5 * C ; C = 1925 / (920*1.5) = 1.3949 ~ 1.40

ROA = PM*TAT ; 13.23% = 11% * F ; F = 13.23%/11% = 1.2027 ~ 1.20

ROE = PM*TAT*EM ; E = 11% * 1.2 *1.33 ; E = 17.556% ~ 17.56%

Final table looks like this:

2017 2018 2019
ROE 18.20% 19.25% 17.56%
EM 1.25 1.4 1.33
ROA 14.60% 13.80% 13.23%
PM 10.00% 9.20% 11.00%
TAT 1.46           1.50 1.2

Comment if you have additonal queries. Please upvote if you like the answer. thanks!